OpenStudy (anonymous):

WILL GIVE MEDAL<<<<<<<<<<<<<<<<<<<<<<<<<<<<< Angle A and B are supplementary. Angle A measures 5x-10 and angle B measures 10x-5. Find the value of x.

OpenStudy (anonymous):

easy measure angle A+measure angleB=180(since the sum of measure of angles of a triangleis 180) 5x-10+10x-5=180 15x-15=180 15x=180+15 15x=195 x=195/15=13 hence x=13

OpenStudy (anonymous):

THANK YOU SO MUCH!